A number when decreased by 25% becomes 90.Find the number
Found 2 solutions by Theo, stanbon:
Answer by Theo(13342)

x is the number.
x - .25x = 90
.75x = 90
x = 90/.75
x = 120
back to your original equation.
120 - .25(120) = 90
120 - 30 = 90
90 = 90
ok
x = 120 is your answer

Answer by stanbon(75887)

A number when decreased by 25% becomes 90.Find the number
Let the number be "x":
Equation:
x - 0.25x = 90
0.75x = 90
(3/4)x = 90
(1/4)x = 30
x = 120
============
Cheers,
Stan H.
============